OpenStudy (anonymous):

A computer is bought for $1380 and sold 10 months later for $520. Assuming the function is linear, find the depreciation formula and the value of the computer after 4 months.

jimthompson5910 (jim_thompson5910):

What's the change in price?

OpenStudy (anonymous):

860?

jimthompson5910 (jim_thompson5910):

over how many months?

OpenStudy (anonymous):

4

jimthompson5910 (jim_thompson5910):

no, read the problem over again

OpenStudy (anonymous):

10? or 6?

jimthompson5910 (jim_thompson5910):

"A computer is bought for $1380 and sold 10 months later for $520"

jimthompson5910 (jim_thompson5910):

so the computer drops by $860 in 10 months Slope = rise/run = -860/10 = -86

jimthompson5910 (jim_thompson5910):

the slope of -86 means the price is dropping by $86 each month

OpenStudy (anonymous):

so is the rise 1380?

jimthompson5910 (jim_thompson5910):

"rise" is the change in price

jimthompson5910 (jim_thompson5910):

change in price = 1380 - 520 = 860

jimthompson5910 (jim_thompson5910):

the change is negative because we have a price drop

jimthompson5910 (jim_thompson5910):

so I guess you should say 520 - 1380 = -860

OpenStudy (anonymous):

so - 860 is after 4 months?

OpenStudy (anonymous):

doesent make sense whats the equation to solve?

jimthompson5910 (jim_thompson5910):

do you see how I'm getting the slope?

jimthompson5910 (jim_thompson5910):

the slope just represents the change in price per month

OpenStudy (anonymous):

yea by subtracting 86 dollors each month

jimthompson5910 (jim_thompson5910):

if the price drops by 860 in 10 months then the price drops by 86 each month

jimthompson5910 (jim_thompson5910):

correct

jimthompson5910 (jim_thompson5910):

now let's say that at time x = 0 months, the price is $1380 this is the y-intercept because x = 0 here

OpenStudy (anonymous):

so is the answer 86 after four months?

jimthompson5910 (jim_thompson5910):

so b = 1380

jimthompson5910 (jim_thompson5910):

slope: m = -86 y-intercept: b = 1380

jimthompson5910 (jim_thompson5910):

y = mx+b y = -86x+b y = -86x+1380

jimthompson5910 (jim_thompson5910):

Plug in x = 10 into y = -86x+1380 and tell me what you get

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

520

jimthompson5910 (jim_thompson5910):

so that equation confirms that in 10 months, the price is $520

jimthompson5910 (jim_thompson5910):

if you were to plug in x = 4, you'd get your answer

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

1036?

jimthompson5910 (jim_thompson5910):

correct

OpenStudy (anonymous):

ok and the depreciation formula is

jimthompson5910 (jim_thompson5910):

you just used the depreciation formula to calculate the price at 4 months

OpenStudy (anonymous):

y= -86x+1380?

jimthompson5910 (jim_thompson5910):

correct

OpenStudy (anonymous):

ok

jimthompson5910 (jim_thompson5910):

-86 tells you how much the price is dropping per month 1380 is the starting price

OpenStudy (anonymous):

ok makes sense

jimthompson5910 (jim_thompson5910):

I'm glad it's clicking

OpenStudy (anonymous):

yea its better this stuff seems hard though

OpenStudy (anonymous):

not very easy to figure out

jimthompson5910 (jim_thompson5910):

but now it should be a bit easier

OpenStudy (anonymous):

yes